A town has a population of 17000 and grows at 4% every year. What will be the population after 12 years, to the nearest whole number?



1
Expert's answer
2021-10-11T10:06:54-0400

P=17000

r=4 %

n=12 years

"A=P(1+\frac{r}{100})^n \\

= 17000(1+0.04)^{12} \\

= 17000 \times 1.04^{12} \\

= 17000 \times 1.601032 \\

= 27217.54 \\

= 27218"